# Probabilistic Feature Augmentation for AIS-Based Multi-Path Long-Term Vessel Trajectory Forecasting

### v5

@@ Results
Discussion
Our study highlights the significance of incorporating a probabilistic model within the trajectory forecasting process to enhance multi-path long-term vessel trajectory forecasting. During training, ground-truth data is used to construct transition probability matrices, providing the model with a nuanced perspective on potential vessel movements without the risk of overfitting. As the model is probabilistically driven, it stores motion statistics without relying heavily on the exact details of training data. This enables the model to distill a set of generalized patterns and rules from the data, which, while retaining a degree of uncertainty, does not compromise the model’s prediction capabilities. This controlled uncertainty enhances its ability to forecast under various conditions and adapt to the complexity of maritime navigation.
Our deep-learning model offers a unique advantage by combining exact coordinates from the Standard Features with conditional probabilities from the Probabilistic Features, allowing for a blend of precision and flexibility in predictive capabilities. This approach enables the model to identify intricate route patterns and make informed decisions that simpler baseline models often overlook. Through the Probabilistic Features, the model learns to navigate a domain where certainty is balanced with possibilities, reflecting real-world conditions where numerous factors can alter a vessel’s trajectory. This dynamic interplay between certainty and uncertainty makes the deep learning model more robust and less likely to fail.
@@ Discussion
Methods
Dataset Description
Our dataset was collected from satellite-based Automatic Identification System (AIS) messages sent by vessels operating in the Gulf of St. Lawrence between 2015 and 2020. The AIS data was acquired from Spire111111 See https://spire.com., and it provides detailed records, including identification number, speed, location, vessel type, and timestamps. After cleaning the tracks, our final dataset consisted of 953 cargo and 391 tanker vessels for training, 51 cargo and 21 tanker vessels for validation, and 252 cargo and 103 tanker vessels for testing. The testing subset represents 20% of the entire dataset, while the validation subset (for deep-learning use) represents 20% of the non-testing data, and the outstanding data constitutes the training dataset.
@@ Methods
Estimation of Route and Destination
Our model for trajectory feature engineering is probabilistic. Its main objective is to predict the vessel’s most likely future location based on the current position, direction data, and historical AIS data. We require the ocean to be segmented into deterministic grids to achieve this. These grids integrate the movement patterns of various vessels over time based on their position data from historical AIS records. Each cell in the grid records when and how frequently vessels have visited that particular cell, storing their motion statistics along with it. The model operates on two levels. Firstly, it predicts the vessel’s route polygon along the trajectory. Secondly, it also predicts the most likely end-grid cell, which represents the completion of the vessel’s journey. Both are represented by their centroids, calculated separately for each AIS message in the vessel trajectory. The step-by-step procedure to find possible routes and destinations for a new trajectory ($\tau_{n}$) is provided in Algorithm 1, which requires as input: $C_{i}$, $S^{C_{i}}_{\tau_{n}}$, and $k$.
@@
TrAISformer Adaptations
As already mentioned, there are only a hand-full of models that are tailored for the long-term trajectory forecasting of AIS data due to the increased uncertainty in working on a large geographical environment over a long temporal range [47], making this problem very challenging to tackle and susceptible to many variables not intrinsically inherent to the positioning data provided by AIS messages. The prior state-of-the-art model [23] in such a task, for instance, was developed to classify instead of regress over data, adopting a space quantization strategy that also helps in reducing uncertainty and achieving better results. Such an approach adopts a sparse representation of AIS data and trains the transformer with multi-head attention. Due to our problem constraints, we were required to adapt their loss function as follows:
@@
The change was necessary because the original implementation of their model failed to converge over our dataset. In such cases, the loss function we adapt measures the variation within the prediction of each feature required by their model. This is a requirement because the authors are using teacher-forcing training, which is also the reason for the increased variability observed in the output results of their model. In our version of the loss function, $\mathbb{L}$ indicates the cross-entropy loss function applied over the quantized model-output results. A last modification for better performance and stability of their model over our dataset refers to masking the embedding using a left-padding strategy instead of a right-padding one to increase the model stability and match the input and output of the different models under testing.
Evaluation Metrics
When evaluating the performance of a probabilistic model, three metrics are used: Precision, Recall, and F1 Score. Precision measures the model’s ability to accurately identify only the relevant instances and is calculated by dividing true positive (TP) predictions by the sum of true positives and false positives (FP). Recall, also known as sensitivity, measures the model’s ability to identify all relevant instances and is calculated by dividing true positives by the sum of true positives and false negatives (FN). F1 Score is a performance assessment metric that combines both Precision and Recall, giving a more comprehensive evaluation of the model’s performance by taking the harmonic mean between these two values. To ensure a fair measure of model performance, each of these metrics is evaluated in a weighted manner.
$\hfill Precision=\frac{TP}{TP+FP}\hfill$ (17)
$Recall=\frac{TP}{TP+FN}\hfill$ (18)
$\hfill F1\ Score=2\times\frac{{Precision\times Recall}}{{Precision+Recall}}\hfill$ (19)
@@
Speed
The speed of a vessel is determined by the ratio of the distance between two points with respect to time. The distance between two AIS messages (points) on the earth’s surface can be determined by the distance between the coordinates of $a,b\in\tau_{i}$ using the great circle distance. The Equation 23 takes into account the latitudes $\mathcal{L}_{at}^{a}$ and $\mathcal{L}_{at}^{b}$, the difference $\bigtriangleup^{\mathcal{L}_{on}}_{ab}$ between the longitudes $\mathcal{L}_{on}^{a}$ and $\mathcal{L}_{on}^{b}$, and the Earth’s radius $\mathbf{R}$ (6,371 km), values are given in radians.
@@
Bearing
The bearing is the direction of a vessel during its trajectory relative to the earth’s surface, calculated as:
@@
Optimization Strategy
The optimization function we use encapsulates the strategy of independently minimizing the Mean Absolute Error (MAE) for both output components, latitude, and longitude, such as formalized below:
@@
Here, $y=\left<\mathcal{L}_{at},\mathcal{L}_{on}\right>$ symbolizes the original coordinates, and $\widehat{y}=\left<\widehat{\mathcal{L}_{at}},\widehat{\mathcal{L}_{on}}\right>$ symbolizes the predicted ones, while $\Theta$ stands as representation of the network parameters. For the optimization, we utilize the Adam optimizer with the learning rate of $1e^{-3}$ and weight decay of $1e^{-4}$ to prevent overfitting.
Model Hyperparameters
To ensure an unbiased evaluation, consistent hyperparameters were utilized across all experiments when constructing the deep learning model. L2 regularization was enforced with a regularization strength of 0.0005. Incremental convolutional blocks were employed with various numbers of filter sizes (256, 256, 128) and kernel sizes (7, 5, 5), with each block being subject to the stipulated L2 regularization. A Dropout Layer was added with a probability rate of 10%. The encoder-decoder architecture utilized bidirectional long- and short-term memory layers with 128 cells. Both the kernel and the recurrent weights in these layers used the ReLU activation function and L2 regularization. A Position-Aware Attention Mechanism was also incorporated to enhance the model’s attention to the input data’s temporal sequence. A time factor of 25% was employed for a balance between self-attention and the capturing of longer-term dependencies within the data sequence. Following the decoder, three dense layers were sequentially stacked. These layers had a decreasing number of neurons (128, 128, 64), and the ReLU function was activated for each.
Computing Environment
Acknowledgments
Author contributions statement
Competing interests
The authors declare no competing interests.

### v6

@@ 5 Results
Legend: Green lines is the input data, Blue the ground-truth, and Red the model forecasting.
@@
6 Conclusion and Remarks
Our study highlights the significance of incorporating a probabilistic model within the trajectory forecasting process to enhance multi-path long-term vessel trajectory forecasting. During training, ground-truth data is used to construct transition probability matrices, providing the deep learning model with a nuanced perspective on potential vessel movements without the risk of overfitting. As the model is probabilistically driven, it stores motion statistics without relying heavily on the exact details of training data. This enables the model to distill a set of generalized patterns and rules from the data, which, while retaining a degree of uncertainty, does not compromise the model’s prediction capabilities. This controlled uncertainty enhances its ability to forecast under various conditions and adapt to the complexity of maritime navigation.
Our deep-learning model offers a unique advantage by combining exact coordinates from the Standard Features with conditional probabilities from the output of probabilistic model (route and endpoint grid cell), allowing for a blend of precision and flexibility in predictive capabilities. This approach enables the model to identify intricate route patterns and make informed decisions that simpler baseline models often overlook. Through the Probabilistic Features, the model learns to navigate a domain where certainty is balanced with possibilities, reflecting real-world conditions where numerous factors can alter a vessel’s trajectory. This dynamic interplay between certainty and uncertainty makes the deep learning model more robust and less likely to fail.
@@ 6 Conclusion and Remarks
Computing Environment
Data Licensing and Training Data Disclosure
The smartWhales initiative and its many research fronts have different license agreements with varying companies and governmental agencies. As a matter of clarification, the research we describe in this paper has used data acquired from Spire under MERIDIAN’s fair-use and non-disclose data license, which is research-intended high-resolution data that is hosted at the Institute for Big Data Analytics (IBDA) from Dalhousie University. Due to licensing agreements, we are not allowed to share the raw data, but we can provide a trained model and means to retrain the models on open-source data from akin maritime regions.
Acknowledgments
Author contributions statement
Competing interests
The authors declare no competing interests.
Appendix A Dataset Description
Our dataset was collected from satellite-based Automatic Identification System (AIS) messages sent by vessels operating in the Gulf of St. Lawrence between 2015 and 2020. The AIS data was acquired from Spire111111 See https://spire.com., and it provides detailed records, including identification number, speed, location, vessel type, and timestamps. After cleaning the tracks, our final dataset consisted of 953 cargo and 391 tanker vessels for training, 51 cargo and 21 tanker vessels for validation, and 252 cargo and 103 tanker vessels for testing. The testing subset represents 20% of the entire dataset, while the validation subset (for deep-learning use) represents 20% of the non-testing data, and the rest constitutes the training dataset.
@@ Appendix A Dataset Description
Appendix B Estimation of Route and Destination
Our probabilistic model’s main objective is to predict the vessel’s most likely future location based on the current position, direction data, and historical AIS data. We require the ocean to be segmented into deterministic grids to achieve this. These grids integrate the movement patterns of various vessels over time based on their position data from historical AIS records. Each cell in the grid records when and how frequently vessels have visited that particular cell, storing their motion statistics along with it. The model operates on two levels. Firstly, it predicts the vessel’s route polygon along the trajectory. Secondly, it also predicts the most likely end-grid cell, which represents the completion of the vessel’s journey. Both are represented by their centroids, calculated separately for each AIS message in the vessel trajectory. The step-by-step procedure to find possible routes and destinations for a new trajectory ($\tau_{n}$) is provided in Algorithm 1, which has inputs: $C_{i}$, $S^{C_{i}}_{\tau_{n}}$, and $k$.
@@ Appendix B Estimation of Route and Destination
Appendix C TrAISformer Adaptations
As already mentioned, there are only a hand-full of models that are tailored for the long-term trajectory forecasting of AIS data due to the increased uncertainty in working on a large geographical environment over a long temporal range [47], making this problem very challenging to tackle and susceptible to many variables not intrinsically inherent to the positioning data provided by AIS messages. The prior state-of-the-art model [23], for instance, was developed to classify instead of regress over data, adopting a space quantization strategy that also helps reduce uncertainty and achieve better results. Such an approach adopts a sparse representation of AIS data and trains the transformer with multi-head attention. Due to problem constraints, we have adapted their loss function as follows:
@@ Appendix C TrAISformer Adaptations
The change was necessary because the original implementation of their model failed to converge over our dataset. In such cases, the loss function we adapt measures the variation within the prediction of each feature required by their model. This is a requirement because the authors are using teacher-forcing training, which is also the reason for the increased variability observed in the output results of their model. In our version of the loss function, $\mathbb{L}$ indicates the cross-entropy loss function applied over the quantized model-output results. A last modification for better performance and stability of their model over our dataset refers to masking the embedding using a left-padding strategy instead of a right-padding, matching the input and output of the different models under testing.
Appendix D Evaluation Metrics
When evaluating the performance of a probabilistic model, three metrics are used: Precision, Recall, and F1 Score. Precision measures the model’s ability to accurately identify only the relevant instances and is calculated by dividing true positive (TP) predictions by the sum of true positives and false positives (FP). Recall, also known as sensitivity, measures the model’s ability to identify all relevant instances and is calculated by dividing true positives by the sum of true positives and false negatives (FN). F1 Score is a performance assessment metric that combines both Precision and Recall, giving a more comprehensive evaluation of the model’s performance by taking the harmonic mean between these two values. To ensure a fair model evaluation, each of the metrics used was evaluated in a weighted manner.
$\hfill\mathcal{P}=\frac{TP}{TP+FP}\hfill$ (17)
$\hfill\mathcal{R}=\frac{TP}{TP+FN}\hfill$ (18)
$\hfill F1\ Score=2\times\frac{{\mathcal{P}\times\mathcal{R}}}{{\mathcal{P}+\mathcal{R}}}\hfill$ (19)
@@ Appendix D Evaluation Metrics
Appendix E Speed
The speed of a vessel is determined by the ratio of the distance between two points with respect to time. The distance between two AIS messages (points) on the earth’s surface can be determined by the distance between the coordinates of $a,b\in\tau_{i}$ using the great circle distance. The Equation 23 uses the latitudes $\mathcal{L}_{at}^{a}$ and $\mathcal{L}_{at}^{b}$, the difference $\bigtriangleup^{\mathcal{L}_{on}}_{ab}$ between the longitudes $\mathcal{L}_{on}^{a}$ and $\mathcal{L}_{on}^{b}$, and the Earth’s radius $\mathbf{R}$ (6,371 km), values are given in radians.
@@ Appendix E Speed
Appendix F Bearing
The bearing is the direction of a vessel relative to the earth’s surface, calculated as:
@@ Appendix F Bearing
Appendix G Optimization Strategy
The optimization function we use encapsulates the strategy of independently minimizing the Mean Absolute Error (MAE) for latitude and longitude, such as formalized below:
@@ Appendix G Optimization Strategy
$y=\left<\mathcal{L}_{at},\mathcal{L}_{on}\right>$ is the original coordinates, and $\widehat{y}=\left<\widehat{\mathcal{L}_{at}},\widehat{\mathcal{L}_{on}}\right>$ is the predicted ones, while $\Theta$ stands as the network parameters. For the optimization process, we utilize the Adam optimizer with the learning rate of $1e^{-3}$ and weight decay of $1e^{-4}$ to prevent overfitting.
Appendix H Model Hyperparameters
Consistent hyperparameters were utilized across all experiments to ensure an unbiased evaluation when constructing the deep learning model. L2 regularization was enforced with a regularization strength of 0.0005. Incremental convolutional blocks were employed with various numbers of filter sizes (256, 256, 128) and kernel sizes (7, 5, 5), with each block being subject to the stipulated L2 regularization. A Dropout Layer was added with a probability rate of 10%. The encoder-decoder architecture utilized bidirectional long- and short-term memory layers with 128 cells. Both the kernel and the recurrent weights in these layers used the ReLU activation function and L2 regularization. A Position-Aware Attention Mechanism was also incorporated to enhance the model’s attention to the input data’s temporal sequence. A time factor of 25% was employed for a balance between self-attention and the capturing of longer-term dependencies within the data sequence. Following the decoder, three dense layers were sequentially stacked. These layers had a decreasing number of neurons (128, 128, 64), and the ReLU function was activated for each.